three times the sum of a number and 3 is 15 less than the number times 7. Find the number.
Found 2 solutions by rfer, macston:
Answer by rfer(16322)

3(x+3)=7x-15
3x+9=7x-15
3x-7x=-15-9
-4x=-24
x=-24/-4
x=6
---------------
3(6+3)=7*6-15
27=27

Answer by macston(5194)

n=a number;
3(n+3)=7n-15
3n+9=7n-15 Subtract 3n from each side.
9=4n-15 Add 15 to each side.
24=4n Divide each side by 4.
6=n ANSWER: The number is 6
CHECK
3(n+3)=7n-15
3(6+3)=7(6)-15
3(9)=42-15
27=27